OpenStudy (anonymous):

How would I solve x-0.1x+0.07(x-0.1x)=69.62

OpenStudy (anonymous):

add and subract all of the coefficients of your x and then divide 69.62 by that coefficient

OpenStudy (jdoe0001):

multiply both sides by 1000, thus you'd end up with \(\bf x-0.1x+0.07(x-0.1x)=69.62\qquad \textit{both sides}\times 1000\\ \quad \\ x-0.1x+0.07(x-0.1x)=69.62\\ \quad \\\implies x-0.1x+0.07x-0.007x=69.62\\ \quad \\ 1000x-100x+70x-7x=69620 \) then solve for "x"

OpenStudy (jdoe0001):

and do as jstew2344 said

OpenStudy (anonymous):

Thanks! :D